Differentiate from the first principle the cube root of x plus cube root of x.
1
Expert's answer
2019-11-27T10:27:09-0500

Given question is to differentiate "x^{1\/3}+x^{1\/3} =2x^{1\/3}"

 Using the first principle

Suppose "f(x)= x^{1\/3}", then we need to evaluate

"lim_{h\to0} \frac{f(x+h)-f(x)}{h}"

Putting the value of f(x) we get,

"lim_{h\to0} \frac{(x+h)^{1/3}-x^{1/3}}{h}"

Multiplying top and bottom by

"(x+h)^{2\/3}+ x^{1\/3}(x+h)^{1\/3}+x^{2\/3}"

Now combine the like terms after multiplying numerator and denominator, many of them will cancel out, we get

"lim_{h\to0} \frac{x+h-x}{h((x+h)^{2/3}+(x+h)^{1/3}x^{1/3}+x^{2/3})}"

Now, cancel x's out on the top:

"lim_{h\to0}\frac{h}{h((x+h)^{2/3}+(x+h)^{1/3}x^{1/3}+x^{2/3})}"

"lim_{h\to0} \frac{1}{(x+h)^{2/3}+(x+h)^{1/3}x^{1/3}+x^{2/3}}"

Now, take the limit, we will get

"\frac{1}{x^{2/3}+x^{2/3}+x^{2/3}} =\frac{1}{3x^{2/3}}"

Thus, the derivative of "f(x) = x^{1\/3}" is "f\prime(x)=\frac{1}{3x^{2/3}}"

So, the derivative of "2x^{1\/3}" will be equal to "2f\prime(x)=\frac{2}{3x^{2/3}}"